OpenStudy (breanneh):

Help I will fan and medal! What is the square root of 142 rounded to the nearest whole number? a. 13 B. 12 c. 11 d. 71

OpenStudy (anonymous):

B 12

OpenStudy (whpalmer4):

11*11 = 121 12*12 = 144 13*13 = 169 71*71 = 5041

OpenStudy (anonymous):

\[\sqrt {142} =11.91 \approx 12\]